lim (x.cosx- sinx / x^2.sinx)

x→0


1
Expert's answer
2021-05-10T13:14:41-0400

"Lim_{x\u21920}\u200b \\frac{xcosx\u2212sinx}{x^2 sinx}\\hspace{1.3cm}(\\frac{0}{0} form)\\newline\u200b\n\\text{By L'Hospital's rule,}\\newline\n\n=Lim_{x\u21920}\u200b \\frac{xcosx\u2212sinx}{x^2 (\\frac{sinx}{x})x}\\newline\n=Lim_{x\u21920}\u200b \\frac{xcosx\u2212sinx}{x^3}\\hspace{1cm}(lim\\frac{sinx}{x}=1)\\newline\n\\text{Differentiate numerator and denominator}\\newline\n=Lim_{x\u21920}\u200b \\frac{-xsinx+cosx\u2212cosx}{3x^2}\\newline\n=Lim_{x\u21920}\u200b \\frac{-xsinx}{3x^2}\\newline\n=Lim_{x\u21920}\u200b \\frac{-x}{3x}(\\frac{sinx}{x})\\newline\n=Lim_{x\u21920}\u200b \\frac{-1}{3}(\\frac{sinx}{x})\\hspace{1cm}(lim\\frac{sinx}{x}=1)\\newline\n=\\frac{-1}{3}"